# Spectral-Spatial Transformer with Active Transfer Learning for Hyperspectral Image Classification

### v1

@@ Abstract
The classification of hyperspectral images (HSI) is a challenging task due to the high spectral dimensionality and limited labeled data typically available for training. In this study, we propose a novel multi-stage active transfer learning (ATL) framework that integrates a Spatial-Spectral Transformer (SST) with an active learning process for efficient HSI classification. Our approach leverages a pre-trained (initially trained) SST model, fine-tuned iteratively on newly acquired labeled samples using an uncertainty-diversity (Spatial-Spectral Neighborhood Diversity) querying mechanism. This mechanism identifies the most informative and diverse samples, thereby optimizing the transfer learning process to reduce both labeling costs and model uncertainty. We further introduce a dynamic freezing strategy, selectively freezing layers of the SST model to minimize computational overhead while maintaining adaptability to spectral variations in new data. One of the key innovations in our work is the self-calibration of spectral and spatial attention weights, achieved through uncertainty-guided active learning. This not only enhances the model’s robustness in handling dynamic and disjoint spectral profiles but also improves generalization across multiple HSI datasets. Additionally, we present a diversity-promoting sampling strategy that ensures the selected samples span distinct spectral regions, preventing overfitting to particular spectral classes. Experiments on benchmark HSI datasets demonstrate that the SST-ATL framework significantly outperforms existing CNN and SST-based methods, offering superior accuracy, efficiency, and computational performance. The source code can be accessed at https://github.com/mahmad000/ATL-SST.
@@ I Introduction
Hyperspectral imaging (HSI) is a powerful remote sensing technique that captures detailed spectral information across narrow wavelength bands, enabling precise identification and classification of diverse land-cover materials [1, 2]. Due to its high spectral resolution, HSI is extensively applied in various fields [3] including remote sensing [4, 5], environmental monitoring [6], food processing [7], red chili adulteration [8, 9], minced meat adulteration [10, 11], forensic [12], biomedical [13, 14], and many more. However, inherent challenges, including high dimensionality, limited labeled samples, and complex spectral-spatial variability, complicate the classification process [15, 16, 17].
Recently, transformer-based architectures have gained prominence in the realm of HSI classification [18, 19]. These architectures leverage self-attention mechanisms that effectively capture long-range dependencies across both spectral and spatial dimensions, significantly enhancing feature representation [20, 21, 22]. Spatial-spectral transformers (SSTs) have emerged as a powerful alternative to convolutional neural networks (CNNs), focusing on modeling interactions between spectral bands and spatial contexts across patches [23, 24, 25]. SSTs exhibit excellent scalability when applied to high-resolution HSIs, efficiently handling large HSI datasets without the necessity for complex pooling operations. This capability has facilitated their successful use in HSI classification, with their flexible architecture contributing to wide-ranging applications. Additionally, SSTs minimize the dependence on manually engineered features by learning hierarchical representations directly from raw pixel data [26], streamlining model development and often leading to enhanced performance. The attention maps produced by SSTs also improve interpretability by identifying the most influential image regions in the model’s predictions, thereby offering valuable insights into the decision-making process [27].
Several studies have explored transformer-based architectures for HSI classification. For example, Shi et al. [28] developed a dual-branch transformer network that effectively captures multi-scale spatial-spectral dependencies by separately processing spectral and spatial tokens. Lan et al. [29] introduced a hybrid-spatial transformer combining convolutional layers for better feature extraction. While these approaches show the effectiveness of transformers, challenges remain [30]. Training large SSTs is computationally demanding due to the quadratic complexity of self-attention, affecting scalability with longer sequences [31]. SSTs may also lack translation invariance compared to CNNs, impacting consistent spatial relationship capture [32]. The tokenization of input images into fixed-size patches can miss fine-grained details, and SSTs often need substantial training data, risking overfitting with smaller datasets [33]. Acquiring labeled HSI data is costly and time-consuming, limiting the feasibility of fully supervised methods [34, 35]. Thus, semi-supervised or active learning (AL) strategies that utilize both labeled and unlabeled data are essential for enhancing classification performance while reducing dependence on annotated samples [36, 37].
AL aims to address the challenges of limited labeled data by selectively querying the most informative samples from a pool of data [38], thereby reducing annotation efforts while maximizing model performance [39, 40]. AL has been effectively utilized in HSI classification to minimize annotation costs. For example, Ahmad et al. [41] implemented fuzziness-based sampling to iteratively update training sets, demonstrating improved performance with fewer labeled samples. Cho et al. [42] proposed an active deep learning framework for HSI classification, integrating CNN with AL to reduce labeling costs and enhance accuracy, with a Markov random field (MRF) refining classification by enforcing label smoothness across classes. Liu et al. [43] introduced MDL4OW, a multitask HSI classification that addresses unknown classes by performing simultaneous classification and reconstruction, identifying unknown classes through reconstruction errors.
Liao et al. [44] presented CGE-AL, a class-wise graph-embedding-based AL framework that utilizes a class-wise graph convolutional network (CGCN) to optimize sample selection for labeling in HSI classification. By measuring uncertainty through graph parameters, CGE-AL iteratively queries and labels the most informative samples, thereby enhancing classification performance with minimal manual annotation. Zhao et al. [45] introduced MAT-ASSAL, an HSI classification framework that combines a multi-attention transformer (MAT) with adaptive superpixel segmentation-based AL to enhance classification performance. MAT captures long-range dependencies and fine-grained local features, while adaptive superpixel segmentation optimizes sample selection for training, preserving essential spatial details. Wang et al. [46] proposed an AL approach that enhances deep classification performance by integrating a co-auxiliary learning strategy and multi-level diversity selection, where an auxiliary network aids in pseudo-labeling and reduces redundancy while ensuring the selection of both representative and uncertain samples. Liu et al. [47] developed an adversarial domain alignment framework with contrastive learning for HSI classification, which enhances cross-domain feature consistency for effective knowledge transfer. By combining spectral and semantic alignment with multiscale feature selection, this method improves robustness and adaptability in few-shot scenarios using jointly optimized adversarial and contrastive losses.
Despite the advancements, deep AL often struggles with the challenges posed by limited labeled samples, which constrain the diversity and quality of selected data, ultimately leading to suboptimal model generalization [48]. Additionally, HSIs present unique challenges due to their intricate spatial-spectral dependencies, which traditional AL methods may not fully capture [49]. These limitations underscore the necessity for an active transfer learning (ATL) framework, wherein an SST can leverage knowledge from labeled source domains to enhance sample selection and feature learning in low-labeled target domains. The integration of AL with transfer learning provides an innovative approach to these challenges, allowing models trained on a source domain to adapt efficiently to a target domain with minimal retraining. This methodology is particularly advantageous for dynamic remote sensing tasks, where the spectral characteristics of new images may differ significantly from those of the training set. By combining SST with ATL, the model can selectively query the most uncertain and diverse samples, iteratively updating the training set and thereby improving the model’s generalization capabilities in the target domain.
In this study, we propose an ATL framework based on an SST (SST-ATL) specifically designed for HSI classification. This framework directly addresses the core challenges of HSI classification, including limited labeled samples, high spectral-spatial variance, and dynamic environmental conditions. It achieves this by utilizing a pre-trained SST model with frozen layers and selectively querying the most valuable samples for annotation. Through this transfer learning paradigm, the pre-trained SST model is fine-tuned on new datasets using AL, allowing the model to adapt to new data distributions while maintaining computational efficiency by querying only the most informative samples. Furthermore, the SST’s self-attention mechanism adeptly captures the complex spatial-spectral dependencies inherent in HSIs, leading to enhanced classification accuracy with minimal annotation effort. The main contributions of this work are as follows:
Figure 1: The workflow of SST with Active Transfer Learning framework for HSI classification. In the top-left, an HSI is processed, followed by patch generation and initial SST training with frozen layers. The active transfer learning loop (top-right) leverages both diversity and uncertainty-based query sampling to iteratively update the training set and pool while reusing the pre-trained SST to minimize computational overhead. The bottom section details the SST encoder, where fused patches are projected through multiple encoding layers. Each patch undergoes multi-head attention, scaled dot-product, and normalization processes. The learned embeddings are enhanced through concatenation and feedforward layers (MLP). This integrated approach aims to refine classification performance with minimal annotation effort.
  1. 1.
    Multi-stage ATL Framework with SST: We introduce a multi-stage ATL framework that integrates SST with an AL process optimized for HSI classification. This framework adopts a pre-trained SST (initially trained and freezer) model in a dynamic setting, where the model is iteratively fine-tuned on newly acquired labeled samples. The novelty lies in the multi-stage adaptation, which not only improves efficiency by minimizing re-training cost but also leverages spectral-spatial context during each stage to effectively handle spectral variability.
  2. 2.
    Self-Calibration of Spectral and Spatial Attention Layers Through Uncertainty-Guided Querying: A key contribution is the uncertainty-guided querying mechanism used to adaptively calibrate the SST’s attention layers. By identifying uncertain samples and incorporating them into the training process, we propose a mechanism that self-calibrates the spectral and spatial attention weights. This self-calibration improves model robustness in environments with dynamic spectral profiles and addresses key limitations in existing SST-based models that struggle to generalize.
  3. 3.
    Diversity-Promoting Sampling Strategy: We propose a novel diversity-promoting (Spatial-Spectral Neighborhood Diversity) sampling strategy that complements uncertainty-based querying in the ATL loop. Unlike existing AL methods that predominantly focus on uncertainty, our approach ensures that selected samples span diverse spectral-spatial regions, thus preventing overfitting to any particular spectral class. This diversity-driven selection is critical for HSI datasets, where different land-cover types may exhibit highly correlated spectral signatures, leading to ambiguous classification boundaries.
  4. 4.
    Dynamic Freezing of SST Layers for Efficient Transfer Learning: We introduce a dynamic layer freezing technique for efficient transfer learning, wherein layers of the SST are selectively frozen and unfrozen based on the incoming data. This method contrasts with static freezing techniques used in transfer learning and achieves a balance between preserving learned spectral-spatial representations and adapting to new spectral variations. This contribution significantly reduces computational overhead, enabling real-time adaptation in resource-constrained environments.
  5. 5.
    Theoretical Formulation and Comprehensive Empirical Validation: We present a rigorous theoretical formulation of the uncertainty-diversity-driven query mechanism and derive bounds for its effectiveness in ATL scenarios. Additionally, we validate our framework on multiple real-world HSI datasets with different spectral characteristics and spatial resolutions, providing a thorough comparison against state-of-the-art methods, including baseline CNNs, and SSTs. Our results demonstrate significant improvements in classification accuracy, robustness, and computational efficiency.
II Proposed Methodology
Given an HSI $\mathbf{X}\in\mathbb{R}^{M\times N\times k}$, where $k$ denotes the number of spectral bands, $M$ is the image height, and $N$ is the width, the model initiates by reshaping $\mathbf{X}$ into a sequence of patches. This is done by dividing $\mathbf{X}$ into overlapping patches of size $W\times W$.
II-A Spectral-Spatial Transformer (SST)
The SST leverages the self-attention mechanism within a transformer architecture to capture spectral and spatial dependencies in HSI. The model has three main components: patch embedding, positional encoding, and the transformer encoder block. The architecture can be formulated as follows.
II-A1 Patch Embedding
Given an input HSI $\mathbf{X}\in\mathbb{R}^{W\times W\times k}$, where $W$ is the spatial window size (patch) and $k$ is the number of spectral bands, we divide the image into patches. The patch embedding operation extracts meaningful features from local regions, producing a set of patches for subsequent processing. Specifically, we use a 3D convolutional layer to extract $p\times p$ patches:
$\mathbf{P}_{i}=\text{Conv3D}(\mathbf{X},\mathbf{W}_{e}),\quad i=1,\dots,N_{p}$ (1)
where $\mathbf{W}_{e}$ are learnable weights of the embedding layer, and $N_{p}$ is the number of patches. Each patch $\mathbf{P}_{i}$ is of dimension $\mathbb{R}^{p\times p\times k}$, and the total number of patches is given by $N_{p}=\left(\frac{W}{p}\right)^{2}$. The embedding dimension is denoted by $d$.
II-A2 Positional Encoding
To incorporate information about the relative positions of the patches, we add a positional encoding $\mathbf{E}_{\text{pos}}$ to the patch embeddings. The positional encoding is computed as follows:
$\mathbf{E}_{\text{pos}}(i,2j)=\sin\left(\frac{i}{10000^{\frac{2j}{d}}}\right),$ (2)
$\mathbf{E}_{\text{pos}}(i,2j+1)=\cos\left(\frac{i}{10000^{\frac{2j}{d}}}\right)$ (3)
where $i$ is the position, $j$ is the index of the dimension, and $d$ is the embedding dimension. The resulting positional encoding is added to the patch embeddings:
$\mathbf{Z}=\mathbf{P}+\mathbf{E}_{\text{pos}}$ (4)
II-A3 Transformer Encoder Block
Each transformer encoder block consists of a multi-head self-attention mechanism followed by a feedforward network (FFN). For each input patch embedding $\mathbf{Z}_{i}$, the self-attention mechanism computes attention scores and aggregates information from all patches. The multi-head attention is defined as:

### v2

# Transformer-Driven Active Transfer Learning for Cross-Hyperspectral Image Classification

@@ Abstract
Hyperspectral image (HSI) classification presents inherent challenges due to high spectral dimensionality, significant domain shifts, and limited availability of labeled data. To address these issues, we propose a novel Active Transfer Learning (ATL) framework built upon a Spatial-Spectral Transformer (SST) backbone. The framework integrates multistage transfer learning with an uncertainty-diversity-driven active learning mechanism that strategically selects highly informative and diverse samples for annotation, thereby significantly reducing labeling costs and mitigating sample redundancy. A dynamic layer freezing strategy is introduced to enhance transferability and computational efficiency, enabling selective adaptation of model layers based on domain shift characteristics. Furthermore, we incorporate a self-calibrated attention mechanism that dynamically refines spatial and spectral weights during adaptation, guided by uncertainty-aware feedback. A diversity-promoting sampling strategy ensures broad spectral coverage among selected samples, preventing overfitting to specific classes. Extensive experiments on benchmark cross-domain HSI datasets demonstrate that the proposed SST-ATL framework achieves superior classification performance compared to conventional approaches. The source code is publicly available at https://github.com/mahmad000/ATL-SST.
@@ I Introduction
Hyperspectral images (HSIs) capture detailed spectral signatures across hundreds of contiguous and narrow wavelength bands, offering rich information that enables accurate classification and identification of diverse materials and land cover types [1]. This fine spectral resolution has established HSIs as a vital tool across numerous domains, including remote sensing [2], environmental monitoring [3], food quality and safety inspection [4, 5, 6], forensic analysis [7], and biomedical imaging [8, 9]. However, despite these widespread applications, achieving robust HSI classification remains a formidable challenge. The high dimensionality of spectral data, the limited availability of labeled samples, and significant spectral-spatial variability due to varying acquisition conditions collectively hinder generalization performance [10, 11]. These issues become particularly pronounced in cross-domain classification tasks, where domain shifts between datasets severely limit model transferability.
In recent years, transformer-based architectures have garnered increasing interest for HSI classification [12, 13, 14], owing to their powerful self-attention mechanisms that capture long-range spectral and spatial dependencies. This capability leads to richer and more discriminative feature representations [15, 16]. Spatial-Spectral Transformers (SSTs), in particular, have emerged as a promising alternative to traditional convolutional neural networks (CNNs). SSTs can model intricate spectral-spatial interactions by operating on image patches, and inherently scale well to high-resolution HSI data [17, 18]. Unlike CNNs, SSTs eliminate the need for handcrafted feature extractors or complex pooling strategies by learning hierarchical representations directly from raw pixels [19]. Furthermore, their interpretability via attention maps enhances trust in model decisions by highlighting salient regions that drive classification outcomes [20, 21].
Building on this foundation, several SST-based models have been proposed better to harness the spectral and spatial properties of HSIs. For instance, Shi et al. [22] introduced a dual-branch transformer architecture that separately captures spectral and spatial dependencies, allowing for more effective multi-scale contextual learning. Nevertheless, deploying SSTs for large-scale HSI tasks presents several challenges. First, the self-attention mechanism incurs quadratic computational complexity with respect to sequence length, posing scalability concerns for longer inputs [23]. Second, unlike CNNs, SSTs lack built-in translation invariance, which may affect the consistency of spatial feature extraction [24]. Third, tokenizing HSIs into fixed-size patches can result in the loss of fine-grained information, particularly in small or irregular structures. Moreover, SSTs typically rely on large amounts of labeled data to achieve strong generalization, making them vulnerable to overfitting under low-label regimes [25]. Given the high cost and effort involved in acquiring labeled HSI samples [26, 27], this reliance on fully supervised learning is often impractical. Consequently, strategies that leverage both labeled and unlabeled data, such as semi-supervised and active learning (AL), have become increasingly important for reducing annotation demands while maintaining classification accuracy [28, 29].
AL, in particular, has gained significant traction in the HSI community as a means of selectively labeling the most informative samples [30, 31]. A number of effective approaches have been introduced to address this challenge [32]. Liu et al. [33] proposed MDL4OW, a multitask framework that combines classification and reconstruction to detect unknown classes via reconstruction errors. Liao et al. [34] introduced CGE-AL, a graph-embedding-based AL method that optimizes sample selection using graph CNNs and uncertainty estimation. Zhao et al. [35] designed MAT-ASSAL, which integrates a multi-attention transformer with adaptive superpixel segmentation to preserve spatial structures during query selection. Similarly, Wang et al. [36] presented a co-auxiliary learning strategy, coupling pseudo-labeling with diverse sample selection to ensure both uncertainty and representativeness. Liu et al. [37] further advanced the field by introducing an adversarial domain alignment framework that incorporates contrastive learning, enhancing robustness and cross-domain generalization under few-shot conditions.
Despite these promising directions, deep AL methods still face key limitations when applied to HSIs. The inherent scarcity of labeled data constrains the diversity of queried samples, often resulting in suboptimal generalization [38]. Additionally, complex spectral-spatial correlations present in HSIs are not always effectively captured by conventional AL techniques [39]. To address these gaps, the research community has started to explore active transfer learning (ATL) frameworks that combine the strengths of AL and transfer learning. In such frameworks, knowledge learned from well-labeled source domains can be transferred to improve sample selection and feature extraction in sparsely labeled target domains. This approach is especially valuable in dynamic remote sensing scenarios, where spectral characteristics can vary substantially over time and location. By enabling domain adaptation with minimal retraining, ATL offers a viable pathway to enhance HSI classification with limited annotation overhead.
Motivated by these insights, this work introduces an SST-based ATL framework (SST-ATL) specifically designed for cross-domain HSI classification. The proposed SST-ATL framework directly tackles critical issues such as limited label availability, spectral-spatial heterogeneity, and domain shifts, providing a comprehensive and adaptive solution. The key contributions of this work are summarized below:
We propose SST-ATL, a unified ATL framework for HSI classification that integrates an SST with a multistage fine-tuning strategy and a dynamic layer freezing mechanism. This design allows for efficient cross-domain adaptation under annotation scarcity by selectively updating model layers based on domain shift characteristics, thereby enhancing generalization while reducing computational overhead. To further improve robustness against spectral variability, a novel self-calibrated attention refinement module is introduced, which dynamically adjusts spatial and spectral attention weights using uncertainty-guided feedback. In addition, we develop a hybrid query strategy that jointly optimizes informativeness, diversity, and spectral coverage through a combined uncertainty-diversity sampling criterion and a spectral diversity-enforcing selection scheme. Extensive evaluations across six standard HSI benchmarks demonstrate that SST-ATL consistently outperforms CNN-, Transformer-, and state-space model-based baselines by +2.5% to +3.8% OA under constrained annotation budgets. The complete implementation is publicly available at https://github.com/mahmad000/ATL-SST.
II Related Works
Recent advancements in HSI classification have explored several paradigms, including transformer-based models, AL, and their integration into ATL. We categorize and review the literature under these themes. Table I summarizes each method’s key contributions, limitations, and the improvements offered by our proposed SST-ATL framework.
Transformer-Based HSI Classification
Transformer-based models, particularly SSTs, have shown exceptional capacity in modeling long-range spectral-spatial dependencies. S2FTNet [40] introduces a dual-branch SST using separate spectral and spatial attention modules, achieving 3–5% gains over CNNs. Zhang et al. [41] propose ELS2T, a resource-efficient SST variant with reduced embedding dimensions. Zhang et al. [42] further design a multi-range SST that attends to spectral-spatial intervals at multiple scales. Ahmad et al. [43] tackle spectral redundancy using a wavelet-based compression (WaveFormer), which reduces FLOPs while preserving spectral detail. Despite these innovations, most transformer-based methods require extensive labeled data and struggle with domain shifts. These limitations motivate the integration of adaptive learning strategies such as AL and ATL to improve generalization under low-label regimes.
TABLE I: Comparative Summary of Recent HSI Classification Approaches.
Method Key Contributions Limitations Proposed SST-ATL
Shi et al. [22] Dual-branch SST for decoupled spectral-spatial attention and multi-scale context modeling Quadratic attention cost, lack of translation invariance, and overfitting risk under low labels Lightweight SST, dynamic freezing for low-label adaptation, and computationally efficient
Zhang et al. [42] Multi-range attention with heads attending to different spectral-spatial scales Fusion overhead and redundant computation across branches Self-calibrated attention mechanism with uncertainty guidance optimizes attention allocation
Ahmad et al. [43] Wavelet-based downsampling before SST to reduce spectral redundancy Optimal wavelet scale selection critical; limited adaptivity to scene heterogeneity Integrates dynamic spectral diversity sampling with reversible compression
Liao et al. [34] Graph embedding-based AL with class-wise sample selection via uncertainty in graph parameters Graph construction is expensive and scales poorly; lacks spectral adaptivity SST-ATL avoids explicit graphs and embeds spectral–spatial diversity in query mechanism
Zhao et al. [35] Multi-attention transformer with adaptive superpixels for active querying Relies heavily on superpixel quality; segmentation biases impact results Patch-free querying driven by calibrated attention and uncertainty
Wang et al. [36] Co-auxiliary learning with pseudo-labeling and uncertainty-based diversity Requires training auxiliary models; complexity grows with iterations No auxiliary network needed; SST-ATL optimizes a unified loss with query feedback
Liu et al. [37] Adversarial domain alignment using contrastive loss to enhance cross-domain generalization Training instability due to adversarial components; contrastive loss tuning is sensitive Domain adaptation via dynamic freezing and progressive fine-tuning improves robustness
Di et al. [44] Dual-module AL with adversarial and inter-class uncertainty sampling Increased training complexity and instability in contrastive pairing under limited labels Simpler and stable sample selection via joint uncertainty–diversity strategy
Zhang et al. [41] Efficient SST (ELS2T) with reduced heads and embeddings for resource-limited settings Reduced capacity on complex datasets; performance drops in high-class overlap SST-ATL retains model expressivity while compressing via reversible transforms
Deng et al. [45] First ATL framework using sparse autoencoders and active querying Based on outdated autoencoder architecture; lacks attention-based adaptability Transformer-based ATL with uncertainty feedback and fine-grained spectral focus
Lin et al. [46] AL-driven transfer learning without prior target knowledge Evaluation limited to three datasets; weak robustness to large domain shifts Evaluated across diverse benchmarks; freezing strategy tailored to shift magnitude
Ahmad et al. [47] Deep AL with 3D CNN and query function using fuzziness, Breaking ties, and MI Computational cost of 3D CNNs high; lacks long-range dependency modeling SST-ATL models long-range features with fewer parameters using hierarchical attention
Yang et al. [48] Conceptual integration of AL and transfer learning for label efficiency High-level design; lacks architectural or query function details SST-ATL provides end-to-end framework with defined transfer, freezing, and query modules
Cao et al. [49] CNN + AL with MRF smoothing for better spatial consistency Lacks global attention modeling and depends on fixed post-processing Self-attention offers adaptive spatial consistency without external smoothers
He et al. [50] Heterogeneous transfer learning from RGB to HSI with mapping and attention Pretrained RGB CNNs underutilize spectral richness; weak generalization under large shifts SST-ATL learns spectral-attentive features directly from HSI with built-in domain adaptation
Active Learning Strategies for HSI
AL has been leveraged to reduce annotation costs by prioritizing informative sample selection. Zhao et al. [35] use adaptive superpixel segmentation with a multi-attention transformer for sample querying. Liao et al. [34] propose class-wise graph-based uncertainty sampling via graph embeddings, though with high computational costs. Wang et al. [36] combine pseudo-labeling with co-auxiliary learning to improve sample representativeness. Di et al. [44] propose ALSN, an adversarially trained Siamese network employing dual uncertainty criteria (AUAL, ICUAL), showing strong label efficiency but with increased model complexity. Cao et al. [49] integrate Markov Random Field smoothing with CNN-based AL to improve spatial coherence, while Ahmad et al. [47] adopt multiple sample selection criteria, including fuzziness, mutual information, and breaking ties in a 3D CNN setting. These methods, while effective, often rely on heuristic sampling or are computationally intensive.
Active Transfer Learning for HSI
The fusion of AL and transfer learning forms ATL frameworks, aimed at learning transferable representations under limited labels. Deng et al. [45] were among the first to propose ATL for HSI using stacked sparse autoencoders. Lin et al. [46] extend this by introducing an AL-driven transfer learning pipeline that does not require prior knowledge of the target domain. Yang et al. [48] provide a conceptual overview of ATL for HSI, but lack architectural detail. He et al. [50] focus on heterogeneous ATL from RGB to HSI using pretrained CNNs, but struggle with spectral misalignment. Our previous work [47] shows that ATL with 3D CNNs can achieve strong results, but at the cost of computational complexity and limited domain adaptation flexibility.
In contrast, our SST-ATL framework offers a principled fusion of SSTs, AL, and ATL by incorporating dynamic layer freezing, uncertainty-calibrated attention, and hybrid uncertainty-diversity querying, providing a robust and computationally efficient solution to cross-domain HSI classification.
Figure 1: Overview of the proposed ATL-SST framework for HSI classification. The model first trains an SST encoder by extracting 3D patch embeddings and learning spatial-spectral representations via multi-head self-attention (MHSA). AL iteratively refines the training set by querying the most informative samples based on uncertainty and diversity criteria, while dynamic layer freezing and self-calibrated attention refinement enhance model generalization and robustness. A fine-tuned SST model is evaluated to achieve efficient and accurate cross-dataset classification with minimal labeled data.
III Proposed Methodology
Given an HSI cube $\mathbf{X}\in\mathbb{R}^{M\times N\times k}$, where $M$ and $N$ are the spatial dimensions and $k$ is the number of spectral bands, we first extract overlapping spatial windows of size $W\times W$ from $\mathbf{X}$. These windows provide localized spatial-spectral context for feature encoding via a transformer-based model. Each window is further divided into smaller subpatches used for embedding.
III-A Spectral-Spectral Transformer (SST)
The SST backbone comprises four stages: patch embedding, positional encoding, self-attention-based encoder blocks, and the final classification head.
Patch Embedding
Each spatial window of size $W\times W\times k$ is processed by a 3D convolutional layer to extract $p\times p$ subpatches. Let $\mathbf{W}_{e}$ be the learnable convolutional kernel bank. The $i$-th embedded subpatch $\mathbf{P}_{i}$ is computed as:
$\mathbf{P}_{i}(u,v,c)=\sum_{m=0}^{p-1}\sum_{n=0}^{p-1}\sum_{q=0}^{k-1}\mathbf{X}(s_{u}+m,s_{v}+n,q)\cdot\mathbf{W}_{e}(m,n,q,c)$ (1)
where $(s_{u},s_{v})$ are the coordinates of the $i$-th subpatch, $p$ is the subpatch size, $c$ indexes output channels, and $N_{p}=\left(\frac{W}{p}\right)^{2}$ is the number of subpatches in a window. These subpatches are projected into a $d$-dimensional feature space.
Positional Encoding
To inject spatial information into the embeddings, a sinusoidal positional encoding $\mathbf{E}_{\text{pos}}$ is added:
$\mathbf{E}_{\text{pos}}(i,2j)$ $=\sin\left(\frac{i}{10000^{2j/d}}\right)$ (2)
$\mathbf{E}_{\text{pos}}(i,2j+1)$ $=\cos\left(\frac{i}{10000^{2j/d}}\right)$ (3)
The encoded patch sequence is $\mathbf{Z}=\mathbf{P}+\mathbf{E}_{\text{pos}}$.
Transformer Encoder Block
Each block applies multi-head self-attention followed by feed-forward processing: